The range of the function f(x)=1-|x-2|, is please solve it hurry up.

Grade:12th pass

1 Answers

Khimraj
6 years ago
Range of (x-2) = (-\infinity∞, ∞)
Range of |x-2| = [0,∞)
Range of -|x-2| = (-∞,0]
So Range of 1-|x-2| = (-∞,1].
Hope it clears. If you like answer then please approve it.